# Hybrid Beamforming Design for RSMA-assisted mmWave Integrated Sensing and Communications

### v1

@@ III-B Proposed PDD-Based Algorithm
The PDD algorithm is a double-loop iterative algorithm. Specifically, the nonconvex problem is transformed into augmented Lagrangian (AL) problems by introducing auxiliary variables and equation constraints, where the inner loop solves the AL subproblem and the outer loop updates the dual variables or the penalty parameters according to the constraint violations. For more details on the PDD algorithm, please refer to [37],[38].
When applying the PDD framework, it is necessary to introduce auxiliary variables and equality constraints to solve the coupling terms and separate multiple constraints in the optimization problem. Next, we stipulate the introduction of auxiliary variables and equality constraints as follows:
1) There can be no coupling terms in all constraints;
2) The same optimization variable cannot appear in more than one constraint;
3) Multiple optimization variables in the same constraint need to be optimized at the same time.
Based on the above regulations, we first introduce a set of auxiliary variables $\boldsymbol{\rm X}$, $\boldsymbol{\rm Y}$, $\{\boldsymbol{\rm Z}_{k}\}$, and $\{\boldsymbol{\rm q}_{k}\}$, $k\in\mathcal{K}$ that satisfy the following equality constraints: $\boldsymbol{\rm X}=\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}$, $\boldsymbol{\rm Y}=\boldsymbol{\rm X}$, $\boldsymbol{\rm Z}_{k}=\boldsymbol{\rm X}$, and $\boldsymbol{\rm q}_{k}=\boldsymbol{\rm c}$. And then, problem P2-1 can be equivalently converted to
[…]
@@
$\boldsymbol{\rm X}=\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D},\boldsymbol{\rm Y}=\boldsymbol{\rm X},$ (29f)
$\boldsymbol{\rm Z}_{k}=\boldsymbol{\rm X},\boldsymbol{\rm q}_{k}=\boldsymbol{\rm c},\forall k\in\mathcal{K},$
where $\mathbb{X}=\{\boldsymbol{\rm F}_{\rm RF},\boldsymbol{\rm F}_{\rm D},\boldsymbol{\rm c},\boldsymbol{\rm u},\boldsymbol{\rm w},\boldsymbol{\rm X},\boldsymbol{\rm Y},\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k}\}$ denotes a set of optimization variables and $\boldsymbol{\rm\alpha}=\left[\alpha_{1},\cdots,\alpha_{K}\right]^{T}$is a vector of user priority weight coefficients. Note that by defining $\boldsymbol{\rm I}_{K+1}=\left[\boldsymbol{\rm e}_{c},\boldsymbol{\rm e}_{1},\cdots,\boldsymbol{\rm e}_{K}\right]$ as a $K+1$ dimensional identity matrix, where $e_{i},i\in\{c\}\cup\mathcal{K}$ denotes the column vector corresponding to matrix $\boldsymbol{\rm I}_{K+1}$, we can further obtain expressions as follows: $\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}=\boldsymbol{\rm X}\boldsymbol{\rm e}_{c}=\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c}$ and $\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}=\boldsymbol{\rm X}\boldsymbol{\rm e}_{i}=\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{i},i\in\mathcal{K}$. And then, the AWMSEs for decoding $s_{k}$ and $s_{c,k}$ are re-described as $\eta_{k}(\boldsymbol{\rm X})$ and $\eta_{c,k}(\boldsymbol{\rm Z}_{k})$, in Eqs. (27) and (28), respectively, at the bottom of this page.
@@
[…]
@@ III-C Proposed Iterative Algorithm for Solving Problem P2-3
In this subsection, we focus on solving the AL problem P2-3. However, the nonconvexity of constraint (29b) causes P2-3 to remain a nonconvex problem. Notice that the constraint (29b) has the structure of the difference-of-convex (DC) function, which can be transformed into a convex constraint by the first-order Taylor expansion, and then the AL problem can be solved iteratively by the concave-convex procedure (CCCP) algorithm. First, constraint (29b) can be rewritten as
@@
where $f_{1}(\boldsymbol{\rm Y})=\gamma_{0}\left(\operatorname{tr}\left(\boldsymbol{\rm Y}^{H}\boldsymbol{\rm\Omega}\boldsymbol{\rm Y}\right)+N_{r}\sigma_{z}^{2}\right)$, $f_{2}(\boldsymbol{\rm Y})=\operatorname{tr}\left(\boldsymbol{\rm Y}^{H}\boldsymbol{\rm\Phi}\boldsymbol{\rm Y}\right)$. At the $t$-th iteration, the first-order Taylor expansion of $f_{2}(\boldsymbol{\rm Y})$ at point $\boldsymbol{\rm Y}^{(t)}$, denoted by $\hat{f}_{2}(\boldsymbol{\rm Y},\boldsymbol{\rm Y}^{(t)})$, can be expressed as
@@
[…]
In the following, we divide the optimized variables of problem P2-4 into five blocks, namely $\{\boldsymbol{\rm w}\}$, $\{\boldsymbol{\rm u}\}$, $\{\boldsymbol{\rm Y},\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k},\boldsymbol{\rm F}_{\rm D}\}$, $\{\boldsymbol{\rm c},\boldsymbol{\rm F}_{\rm RF}\}$, and $\{\boldsymbol{\rm X}\}$. A parallel optimization strategy is used for the variables within the blocks, which reduces computational resources and time, and then iteratively updates between the blocks based on the BCD algorithm. The specific optimization steps for variable blocks are given as follows.
$\bm{Step1:}$ we optimize the equalizer vector $\{\boldsymbol{\rm w}\}$ by fixing other variables. Based on the derivation in Section III.A, the optimal equalizers, denoted by $w_{c,k}^{opt}$ and $w_{k}^{opt}$, $k\in\mathcal{K}$, can be updated as
$w_{c,k}^{opt}$ $=(\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c})^{H}\boldsymbol{\rm h}_{k}(T_{c,k}(\boldsymbol{\rm Z}_{k}))^{-1},$ (35a)
$w_{k}^{opt}$ $=(\boldsymbol{\rm X}\boldsymbol{\rm e}_{k})^{H}\boldsymbol{\rm h}_{k}(T_{k}(\boldsymbol{\rm X}))^{-1},$ (35b)
where $T_{c,k}(\boldsymbol{\rm Z}_{k})$ and $T_{k}(\boldsymbol{\rm X})$ are obtained by taking $\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}=\boldsymbol{\rm X}=\boldsymbol{\rm Z}_{k}$ into $T_{c,k}$ and $T_{k}$, respectively.
$\bm{Step2:}$ we optimize the weight vector $\{\boldsymbol{\rm u}\}$ by fixing other variables. First, substitute $\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}=\boldsymbol{\rm X}=\boldsymbol{\rm Z}_{k}$ into Eqs. (20) and (21) to obtain $e_{c,k}^{\rm MMSE}(\boldsymbol{\rm Z}_{k})$ and $e_{k}^{\rm MMSE}(\boldsymbol{\rm X})$, respectively. Then the optimal weights, denoted by $u_{c,k}^{opt}$ and $u_{k}^{opt}$, $k\in\mathcal{K}$, are given by
$u_{c,k}^{opt}=(e_{c,k}^{\rm MMSE}(\boldsymbol{\rm Z}_{k})\ln 2)^{-1},u_{k}^{opt}=(e_{k}^{\rm MMSE}(\boldsymbol{\rm X})\ln 2)^{-1}.$ (36)
$\bm{Step3:}$ by fixing the other variables, we optimize the block variables $\{\boldsymbol{\rm Y},\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k},\boldsymbol{\rm F}_{\rm D}\}$ in parallel. Specifically, it can be decomposed into three optimization problems. First, the subproblem for variable $\boldsymbol{\rm Y}$ can be expressed as
@@
To obtain the optimal solution of problem (37), we present the following proposition.
Proposition 2
The optimal solution of problem (37), denoted by $\boldsymbol{\rm Y}^{opt}(v_{1})$, is expressed as
@@
where $v_{1}\geq 0$ is the Lagrange multiplier that needs to satisfy the KKT condition of problem (37), which can be obtained by one-dimensional line search.
[…]
Note that in the $(t+1)$ iteration, $\boldsymbol{\rm Y}^{(t)}$ is first updated by $\boldsymbol{\rm Y}^{(t)}=\boldsymbol{\rm Y}^{opt}$, and then $\boldsymbol{\rm Y}^{(t)}$ is taken into Eq. (38) to obtain the new $\boldsymbol{\rm Y}^{opt}$, i.e., updated in a recursive method.
Since variables $\boldsymbol{\rm Z}_{k}$ and $\boldsymbol{\rm q}_{k}$ are under the same constraint (29c), they need to be optimized simultaneously. By fixing the other variables in problem P2-4, the subproblem for variables $\boldsymbol{\rm Z}_{k}$ and $\boldsymbol{\rm Z}_{k}$ can be given by
$\min_{\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k}}$ $\|\boldsymbol{\rm X}-\boldsymbol{\rm Z}_{k}+\rho\boldsymbol{\rm\Lambda}_{2,k}\|_{F}^{2}+\left\|\boldsymbol{\rm c}-\boldsymbol{\rm q}_{k}+\rho\boldsymbol{\rm\lambda}_{c,k}\right\|_{F}^{2}$ (39a)
s.t. $u_{c,k}\left|w_{c,k}\right|^{2}\left(\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c}\right|^{2}+\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{i}\right|^{2}\right)$
$-2u_{c,k}\mathfrak{R}\left\{w_{c,k}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c}\right\}+\boldsymbol{\rm 1}^{T}\boldsymbol{\rm q}_{k}\leq\varpi_{k},$ (39b)
where $\varpi_{k}=g-u_{c,k}\left|w_{c,k}\right|^{2}\sigma_{k}^{2}-u_{c,k}+\log_{2}\left(u_{c,k}\right)$ denotes a constant. The optimal solution of problem (39) can be obtained by the KKT condition. We also present the proposition as follows:
Proposition 3
In problem (39), the optimal solutions of $\boldsymbol{\rm Z}_{k}$ and $\boldsymbol{\rm q}_{k}$, denoted by $\boldsymbol{\rm Z}_{k}^{opt}(v_{2,k})$ and $\boldsymbol{\rm q}_{k}^{opt}(v_{2,k})$, are given by
$\boldsymbol{\rm Z}_{k}^{opt}(v_{2,k})=$ $\left(\boldsymbol{\rm I}_{N_{t}}+v_{2,k}u_{c,k}\left|w_{c,k}\right|^{2}\boldsymbol{\rm h}_{k}\boldsymbol{\rm h}_{k}^{H}\right)^{-1}\bigg{(}\boldsymbol{\rm X}+\rho\boldsymbol{\rm\Lambda}_{2,k}$
$+v_{2,k}u_{c,k}w_{c,k}^{*}\boldsymbol{\rm h}_{k}\boldsymbol{\rm e}_{c}^{T}\bigg{)},$ (40)
$\boldsymbol{\rm q}_{k}^{opt}(v_{2,k})=$ $\boldsymbol{\rm c}+\rho\boldsymbol{\rm\lambda}_{c,k}-\frac{v_{2,k}}{2}\boldsymbol{\rm 1},$ (41)
where $v_{2,k}\geq 0$ denotes the Lagrange multiplier which can be obtained by the bisection method.
[…]
@@
By taking the first-order derivative of the variable $\boldsymbol{\rm F}_{\rm D}$ for the above subproblem (42), we obtain the optimal solution, denoted by $\boldsymbol{\rm F}_{\rm D}^{opt}$, which can be given by
$\boldsymbol{\rm F}_{\rm D}^{opt}=\boldsymbol{\rm F}_{\rm RF}^{\dagger}\left(\boldsymbol{\rm X}+\rho\boldsymbol{\rm\Lambda}_{3}\right).$ (43)
$\bm{Step4:}$ we optimize the block variables $\{\boldsymbol{\rm c},\boldsymbol{\rm F}_{\rm RF}\}$ by fixing other variables. The subproblem for the variable $\boldsymbol{\rm c}$ is given by
@@
The problem (44) is a convex problem that can be easily optimized by the KKT condition. The optimal solution for the $i$ th element of the vector $\boldsymbol{\rm c}$, denoted by $C_{i}^{opt},i\in\mathcal{K}$, has the closed form as follows:
@@
This subproblem is a non-convex quadratic programming problem with unit modulus constraints. In order to solve this problem, we first transform it into a tractable form, as follows:
@@
where $\boldsymbol{\rm C}=\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}$, $\boldsymbol{\rm B}=\left(\boldsymbol{\rm X}+\rho\boldsymbol{\rm\Lambda}_{3}\right)\boldsymbol{\rm F}_{\rm D}^{H}$. And then, we apply the single-iteration BCD algorithm shown in [43, Appendix B] to recursively solve problem (47). The core of the algorithm lies in iteratively updating each entry in the variable $\boldsymbol{\rm F}_{\rm RF}$. The reader is referred to [43, Appendix B] for more details.
$\bm{Step5:}$ we optimize the variable $\boldsymbol{\rm X}$ by fixing other variables. Then the specific subproblem can be expressed as
@@
Problem (48) is a quadratic convex programming problem, and by introducing the Lagrange multiplier $v_{3}\geq 0$, we can obtain the optimal $\boldsymbol{\rm X}^{opt}(v_{3})$, which is given by the following proposition:
Proposition 4
The vectorization of the optimal solution to problem (48), denoted by $\boldsymbol{\rm x}^{opt}(v_{3})$, can be expressed as
$\boldsymbol{\rm x}^{opt}(v_{3})=\operatorname{vec}\left(\boldsymbol{\rm X}^{opt}(v_{3})\right)=\bigg{(}\bigg{(}\frac{K+2}{2\rho}+v_{3}\bigg{)}\boldsymbol{\rm I}_{N_{t}\times(K+1)}+$
$\bigg{(}\sum_{i\in\mathcal{K}}\boldsymbol{\rm e}_{i}\boldsymbol{\rm e}_{i}^{H}\bigg{)}^{T}\otimes\bigg{(}\sum_{k\in\mathcal{K}}\alpha_{k}u_{k}\left|w_{k}\right|^{2}\boldsymbol{\rm h}_{k}\boldsymbol{\rm h}_{k}^{H}\bigg{)}\bigg{)}^{-1}\operatorname{vec}(\boldsymbol{\rm V}),$ (49)
where
$\boldsymbol{\rm V}=$ $\sum_{k\in\mathcal{K}}\alpha_{k}u_{k}w_{k}^{*}\boldsymbol{\rm h}_{k}\boldsymbol{\rm e}_{k}^{H}+\frac{1}{2\rho}\bigg{(}\boldsymbol{\rm Y}+\sum_{k\in\mathcal{K}}\boldsymbol{\rm Z}_{k}+\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}$
$-\rho\bigg{(}\boldsymbol{\rm\Lambda}_{1}+\sum_{k\in\mathcal{K}}\boldsymbol{\rm\Lambda}_{2,k}+\boldsymbol{\rm\Lambda}_{3}\bigg{)}\bigg{)}.$ (50)
The Lagrange multiplier $v_{3}$ can be obtained by the bisection method.

### v2

@@ III-B Proposed PDD-Based Algorithm
In the following, we propose the PDD-based algorithm to solve problem P2-1. We first introduce a set of auxiliary variables $\boldsymbol{\rm X}$, $\boldsymbol{\rm Y}$, $\{\boldsymbol{\rm Z}_{k}\}$, and $\{\boldsymbol{\rm q}_{k}\}$, $k\in\mathcal{K}$ that satisfy the following equality constraints: $\boldsymbol{\rm X}=\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}$, $\boldsymbol{\rm Y}=\boldsymbol{\rm X}$, $\boldsymbol{\rm Z}_{k}=\boldsymbol{\rm X}$, and $\boldsymbol{\rm q}_{k}=\boldsymbol{\rm c}$. And then, problem P2-1 can be equivalently converted to
$\textbf{{{P}}2-2:}\min_{\mathbb{X}}$ $\sum_{k\in\mathcal{K}}\alpha_{k}\eta_{k}(\boldsymbol{\rm X})-\boldsymbol{\rm\alpha}^{T}\boldsymbol{\rm c}$ (11a)
@@
$\boldsymbol{\rm X}=\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D},\boldsymbol{\rm Y}=\boldsymbol{\rm X},\boldsymbol{\rm Z}_{k}=\boldsymbol{\rm X},\boldsymbol{\rm q}_{k}=\boldsymbol{\rm c},$ (11f)
where $\mathbb{X}=\{\boldsymbol{\rm F}_{\rm RF},\boldsymbol{\rm F}_{\rm D},\boldsymbol{\rm c},\boldsymbol{\rm u},\boldsymbol{\rm w},\boldsymbol{\rm X},\boldsymbol{\rm Y},\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k}\}$ denotes a set of optimization variables and $\boldsymbol{\rm\alpha}=\left[\alpha_{1},\cdots,\alpha_{K}\right]^{T}$is a vector of user priority weight coefficients.
@@
$\textbf{{{P}}2-3:}\min_{\mathbb{X}}$ $\sum_{k\in\mathcal{K}}\alpha_{k}\eta_{k}(\boldsymbol{\rm X})-\boldsymbol{\rm\alpha}^{T}\boldsymbol{\rm c}+g_{\rho}(\mathbb{X})$ (12a)
@@ III-C Proposed Iterative Algorithm for Solving Problem P2-3
According to the concave-convex procedure (CCCP) algorithm, the constraint (13b) can be transformed as
@@
where $f_{1}(\boldsymbol{\rm Y})=\gamma_{0}\left(\operatorname{tr}\left(\boldsymbol{\rm Y}^{H}\boldsymbol{\rm\Omega}\boldsymbol{\rm Y}\right)+N_{r}\sigma_{z}^{2}\right)$ and $f_{2}(\boldsymbol{\rm Y})=\operatorname{tr}\left(\boldsymbol{\rm Y}^{H}\boldsymbol{\rm\Phi}\boldsymbol{\rm Y}\right)$. At the $t$-th iteration, the first-order Taylor expansion of $f_{2}(\boldsymbol{\rm Y})$ at point $\boldsymbol{\rm Y}^{(t)}$, denoted by $\hat{f}_{2}(\boldsymbol{\rm Y},\boldsymbol{\rm Y}^{(t)})$, can be expressed as
@@
$\textbf{{{P}}2-4:}\min_{\mathbb{X}}$ $\sum_{k\in\mathcal{K}}\alpha_{k}\eta_{k}(\boldsymbol{\rm X})-\boldsymbol{\rm\alpha}^{T}\boldsymbol{\rm c}+g_{\rho}(\mathbb{X})$ (16a)
@@
In the following, we divide the optimized variables of problem P2-4 into five blocks, namely $\{\boldsymbol{\rm w}\}$, $\{\boldsymbol{\rm u}\}$, $\{\boldsymbol{\rm Y},\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k},\boldsymbol{\rm F}_{\rm D}\}$, $\{\boldsymbol{\rm c},\boldsymbol{\rm F}_{\rm RF}\}$, and $\{\boldsymbol{\rm X}\}$. We define the accuracy tolerance $\epsilon^{in}$ and the maximum number of iterations $N_{max}^{in}$, initialize the iteration number $t=0$, and repeat the steps until the difference between the OF values of two adjacent iterations is less than $\epsilon^{in}$ or $t\geq N_{max}^{in}$. The specific optimization steps for variable blocks are given as follows.
$\bm{Step1}$: Optimize the equalizer vector $\{\boldsymbol{\rm w}\}$ by fixing other variables. Based on the derivation in Section III.A, the optimal equalizers can be updated as $w_{c,k}^{opt}=(\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c})^{H}\boldsymbol{\rm h}_{k}(T_{c,k}(\boldsymbol{\rm Z}_{k}))^{-1}$ and $w_{k}^{opt}=(\boldsymbol{\rm X}\boldsymbol{\rm e}_{k})^{H}\boldsymbol{\rm h}_{k}(T_{k}(\boldsymbol{\rm X}))^{-1}$, where $T_{c,k}(\boldsymbol{\rm Z}_{k})$ and $T_{k}(\boldsymbol{\rm X})$ are obtained by taking $\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}=\boldsymbol{\rm X}=\boldsymbol{\rm Z}_{k}$ into $T_{c,k}$ and $T_{k}$, respectively.
$\bm{Step2}$: Optimize the weight vector $\{\boldsymbol{\rm u}\}$ by fixing other variables. First, substitute $\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}=\boldsymbol{\rm X}=\boldsymbol{\rm Z}_{k}$ into Eqs. (9) to obtain $e_{c,k}^{\rm MMSE}(\boldsymbol{\rm Z}_{k})$ and $e_{k}^{\rm MMSE}(\boldsymbol{\rm X})$, respectively. Then, the optimal weights, denoted by $u_{c,k}^{opt}$ and $u_{k}^{opt}$, are given by $u_{c,k}^{opt}=(e_{c,k}^{\rm MMSE}(\boldsymbol{\rm Z}_{k})\ln 2)^{-1}$ and $u_{k}^{opt}=(e_{k}^{\rm MMSE}(\boldsymbol{\rm X})\ln 2)^{-1}$.
$\bm{Step3}$: By fixing the other variables, optimize the block variables $\{\boldsymbol{\rm Y},\boldsymbol{\rm Z}_{k},\boldsymbol{\rm q}_{k},\boldsymbol{\rm F}_{\rm D}\}$. First, the subproblem for variable $\boldsymbol{\rm Y}$ can be expressed as
@@
The optimal solution of problem (15), denoted by $\boldsymbol{\rm Y}^{opt}(v_{1})$, can be obtained by one-dimensional line search, as follows:
@@
where $v_{1}\geq 0$ is the Lagrange multiplier that needs to satisfy the KKT condition. Note that in the $(t+1)$ iteration, $\boldsymbol{\rm Y}^{(t)}$ is first updated by $\boldsymbol{\rm Y}^{(t)}=\boldsymbol{\rm Y}^{opt}$, and then $\boldsymbol{\rm Y}^{(t)}$ is taken into Eq. (18) to obtain the new $\boldsymbol{\rm Y}^{opt}$, i.e., updated in a recursive method.
Similarly, the optimal solutions of $\boldsymbol{\rm Z}_{k}$ and $\boldsymbol{\rm q}_{k}$, denoted by $\boldsymbol{\rm Z}_{k}^{opt}(v_{2,k})$ and $\boldsymbol{\rm q}_{k}^{opt}(v_{2,k})$, are given by $\boldsymbol{\rm Z}_{k}^{opt}(v_{2,k})=\big{(}\boldsymbol{\rm I}_{N_{t}}+v_{2,k}u_{c,k}\left|w_{c,k}\right|^{2}\boldsymbol{\rm h}_{k}\boldsymbol{\rm h}_{k}^{H}\big{)}^{-1}\big{(}\boldsymbol{\rm X}+\rho\boldsymbol{\rm\Lambda}_{2,k}+v_{2,k}u_{c,k}w_{c,k}^{*}\boldsymbol{\rm h}_{k}\boldsymbol{\rm e}_{c}^{T}\big{)}$ and $\boldsymbol{\rm q}_{k}^{opt}(v_{2,k})=\boldsymbol{\rm c}+\rho\boldsymbol{\rm\lambda}_{c,k}-\frac{v_{2,k}}{2}\boldsymbol{\rm 1}$, where $v_{2,k}\geq 0$ denotes the Lagrange multiplier which can be obtained by the bisection method.
@@
By taking the first-order derivative of the variable $\boldsymbol{\rm F}_{\rm D}$ for the above subproblem (19), we obtain the optimal solution $\boldsymbol{\rm F}_{\rm D}^{opt}=\boldsymbol{\rm F}_{\rm RF}^{\dagger}\big{(}\boldsymbol{\rm X}+\rho\boldsymbol{\rm\Lambda}_{3}\big{)}$.
$\bm{Step4}$: Optimize the block variables $\{\boldsymbol{\rm c},\boldsymbol{\rm F}_{\rm RF}\}$ by fixing other variables. The subproblem for the variable $\boldsymbol{\rm c}$ is given by
@@
The problem (20) is a convex problem that can be easily optimized by the Karush-Kuhn-Tucker (KKT) condition. The optimal solution for the $i$-th element of the vector $\boldsymbol{\rm c}$, denoted by $C_{i}^{opt},i\in\mathcal{K}$, has the closed form as follows:
@@
To solve this non-convex quadratic programming problem, we transform it into a tractable form, as follows:
@@
where $\boldsymbol{\rm C}=\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}$, $\boldsymbol{\rm B}=\left(\boldsymbol{\rm X}+\rho\boldsymbol{\rm\Lambda}_{3}\right)\boldsymbol{\rm F}_{\rm D}^{H}$.
$\bm{Step5}$: Optimize the variable $\boldsymbol{\rm X}$ by fixing other variables. Then the specific subproblem can be expressed as
@@
Problem (24) is a quadratic convex programming problem, and by introducing the Lagrange multiplier $v_{3}\geq 0$, we can obtain the optimal $\boldsymbol{\rm X}^{opt}(v_{3})$, which is given by $\boldsymbol{\rm x}^{opt}(v_{3})\!=\!\operatorname{vec}\left(\boldsymbol{\rm X}^{opt}(v_{3})\right)\!=\!\big{(}(\frac{K+2}{2\rho}\!+\!v_{3})\boldsymbol{\rm I}_{N_{t}\times(K+1)}\!+\!(\sum_{i\in\mathcal{K}}\boldsymbol{\rm e}_{i}\boldsymbol{\rm e}_{i}^{H})^{T}\!\otimes\!(\sum_{k\in\mathcal{K}}\alpha_{k}u_{k}\left|w_{k}\right|^{2}\boldsymbol{\rm h}_{k}\boldsymbol{\rm h}_{k}^{H})\big{)}^{\!-\!1}\operatorname{vec}(\boldsymbol{\rm V})$, where $\boldsymbol{\rm V}=\sum_{k\in\mathcal{K}}\alpha_{k}u_{k}w_{k}^{*}\boldsymbol{\rm h}_{k}\boldsymbol{\rm e}_{k}^{H}+\frac{1}{2\rho}\big{(}\boldsymbol{\rm Y}+\sum_{k\in\mathcal{K}}\boldsymbol{\rm Z}_{k}+\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}-\rho(\boldsymbol{\rm\Lambda}_{1}+\sum_{k\in\mathcal{K}}\boldsymbol{\rm\Lambda}_{2,k}+\boldsymbol{\rm\Lambda}_{3})\big{)}$, and $\otimes$ denotes the Kronecker products between two matrices. The Lagrange multiplier $v_{3}$ can be obtained by the bisection method.
III-D Summary of the Proposed WMMSE-PDD Algorithm
After each inner iteration is completed, we need to update the Lagrange multipliers and the penalty parameter of the outer loop. We first define the constraint violation in the $m$-th outer iteration, denoted by $h\left(\mathbb{X}^{(m)}\right)$, expressed by $h\left(\mathbb{X}^{(m)}\right)\!=\!\max\{\|\boldsymbol{\rm X}\!-\!\boldsymbol{\rm Y}\|_{F},\|\boldsymbol{\rm X}\!-\!\boldsymbol{\rm Z}_{k}\|_{F},\|\boldsymbol{\rm c}\!-\!\boldsymbol{\rm q}_{k}\|_{2},\|\boldsymbol{\rm X}\!-\!\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\|_{F}\}$. In the $(m+1)$-th outer iteration, we introduce the variable $\zeta^{(m+1)}$ associated with the constraint violation $h(\mathbb{X}^{(m)})$, usually set to $\zeta^{(m+1)}=0.9h(\mathbb{X}^{(m)})$, and then give the update rule as follows: when $h(\mathbb{X}^{(m+1)})\geq\zeta^{(m+1)}$, the penalty parameter can updated according to $\rho^{(m+1)}=c_{\rho}\rho^{(m)}(0<c_{\rho}<1)$; otherwise, the Lagrange multipliers need to be updated, which can be given by
$\begin{cases}&\hskip-11.38092pt\boldsymbol{\rm\Lambda}_{1}^{(m+1)}\!=\!\boldsymbol{\rm\Lambda}_{1}^{(m)}\!+\!\frac{\boldsymbol{\rm X}^{(m)}-\boldsymbol{\rm Y}^{(m)}}{\rho^{(m)}},\boldsymbol{\rm\Lambda}_{2,k}^{(m+1)}\!=\!\boldsymbol{\rm\Lambda}_{2,k}^{(m)}\!+\!\frac{\boldsymbol{\rm X}^{(m)}\!-\!\boldsymbol{\rm Z}_{k}^{(m)}}{\rho^{(m)}};\\ &\hskip-11.38092pt\boldsymbol{\rm\Lambda}_{3}^{(m+1)}\!\!=\!\!\boldsymbol{\rm\Lambda}_{3}^{(m)}\!+\!\frac{\boldsymbol{\rm X}^{(m)}\!-\!\boldsymbol{\rm F}_{\rm RF}^{(m)}\boldsymbol{\rm F}_{\rm D}^{(m)}}{\rho^{(m)}},\boldsymbol{\rm\lambda}_{c,k}^{(m+1)}\!=\!\boldsymbol{\rm\lambda}_{c,k}^{(m)}\!\!+\!\!\frac{\boldsymbol{\rm c}^{(m)}\!-\!\boldsymbol{\rm q}_{k}^{(m)}}{\rho^{(m)}}.\end{cases}$ (25)